OpenStudy (anonymous):

Choose the correct simplification of (4x3 + 3x2 − 6x) − (10x3 + 3x2). 6x3 + 6x −6x3 − 6x2 − 6x −6x3 − 6x 6x3 + 6x2 + 6x

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I know its between B and C because it starts with -6x3 since 4x3 - 10x3 is -6x3

OpenStudy (anonymous):

all you have to do is distribute the negative to the expressions in the 2nd parenthese and then add it up

OpenStudy (anonymous):

btw one of your choices is the correct one

OpenStudy (mathstudent55):

The fisrt set of parentheses is unnecessary, so just drop them. Outside the second set of parentheses, there is a negative sign. That is going to change every sign inside the parentheses. Rewrite this by omitting the first set of parentheses and change every sign inside the second set of parentheses. Then you simplify by combining like terms.

OpenStudy (anonymous):

3x2 + 3x2 = 6x2 but why is it going to be - 3x2 ?

OpenStudy (mathstudent55):

Read my comments above.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

what's a positive times a negative?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

well whats 3x^2-3x^2

OpenStudy (anonymous):

0

OpenStudy (anonymous):

o so its C since u ake that out

OpenStudy (anonymous):

take*

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

(4x3 + 3x2 − 6x) − (10x3 + 3x2) = 4x3 + 3x2 − 6x − 10x3 - 3x2 = 4x3 - 10x3 + 3x2 - 3x2 - 6x = -6x3 - 6x